### v1

@@ Abstract
Solar and wind power are cost-competitive with fossil fuels, yet their intermittent nature presents challenges. Significant temporal and geographic differences in land, wind, and solar resources suggest that long-distance transmission could be particularly beneficial. Using a detailed, open-source model, we analyze optimal transmission expansion jointly with storage, generation, and hourly operations across the three primary interconnects in the United States. Transmission expansion offers far more benefits in a high-renewable system than in a system with mostly conventional generation. Yet while an optimal nationwide plan would have more than triple current interregional transmission, transmission decreases the cost of a 100% clean system by only 4% compared to a plan that relies solely on current transmission. Expanding capacity only within existing interconnects can achieve most of these savings. Adjustments to energy storage and generation mix can leverage the current interregional transmission infrastructure to build a clean power system at a reasonable cost.
@@ 1 Introduction
Although solar and wind power are affordable and likely central to efforts to decarbonize the world economy, their variability presents a considerable challenge to existing systems dominated by easily controllable thermal power plants. Interregional transmission can help smooth imbalances in supply and demand over space while transferring power from places relatively rich in solar and wind resources to places with poorer options. For example, compared to other regions, the desert Southwest receives more solar radiation, while the Great Plains have abundant wind resources. These aspects of solar and wind contrast sharply with existing generation. Thermal power plants perform nearly the same wherever they are placed. It is unclear how suitable existing transmission lines are for the emerging energy system with different interregional and intertemporal gains from trade. At the same time, extreme weather events and reliability impacts (e.g., heat waves and wildfires in California and winter storms in Texas) have increased discussion of the need for strong connections between U.S. grids.
Existing interregional transmission capacity was not built with a high-renewable future in mind, and building new transmission lines is costly economically, logistically, environmentally, and politically. Historically, building new transmission lines has often been fraught with controversies about where to site transmission lines, how to finance them, and how costs should be allocated to buyers, sellers, and electricity customers. While enhanced trade between regions generally benefits all sides, individual stakeholders, such as existing transmission line owners and locally advantaged existing generators, can lose rents amid enhanced competition (Hausman, 2024). New transmission lines might also pass through regions that do not benefit from them, and these communities may understandably resist their siting, especially when transmission line construction has considerable environmental and biodiversity impacts (Biasotto and Kindel, 2018; Marshall and Baxter, 2002; Söderman, 2006; Hyde, Bohlman and Valle, 2018). In addition, ERCOT, the Texas-only interconnect, is especially isolated and likely to resist connections with other regions given its unique market structure and aversion to oversight from the Federal Energy Regulatory Commission (FERC). These challenges cast doubt on the likelihood that interregional transmission will expand in line with conventional measures of benefits and costs.
When these challenges are paired with a perception that transmission expansion is an essential part of the decarbonization puzzle (Brown and Botterud, 2021; Pacala et al., 2021; Moch and Lee, 2022; Joskow, 2020; Davis, Hausman and Rose, 2023), it highlights the importance of evaluating the benefits and costs of interregional transmission upgrades, taking into account the full portfolio of alternatives, both within and between the existing interconnects. If future transmission expansion is uncertain, some may wonder whether near-term investments in solar, wind, and other infrastructure will be compromised if transmission expansion does not proceed as expected. While there is a significant body of literature addressing the optimal design of electricity systems for the U.S. (Brown et al., 2018; Mai et al., 2018; Jayadev, Leibowicz and Kutanoglu, 2020), few studies consider 100% decarbonization. Even fewer studies have evaluated the value of transmission in a 100% decarbonization future. Given the scale and useful lifetimes of these investments, it is important to consider the rapidly falling costs of storage and other technologies that could serve as substitutes.
Findings from existing work about transmission are divergent. Using a modified version of OSeMOSYS (Howells et al., 2011), (Jayadev, Leibowicz and Kutanoglu, 2020) illustrate the limited value of long-distance transmission in the least cost scenario with no carbon policy applied. In contrast, the recent study by (Brown and Botterud, 2021) finds that interregional transmission upgrades in a zero-carbon electricity system can reduce the average cost of decarbonized power from over $100 per MWh to about $70 per MWh. However, this measure compares ideal “greenfield" transmission systems and pays no attention to existing transmission capacities. Moreover, it does not consider seasonal balancing options, such as green hydrogen, nuclear power or carbon capture and storage (CCS) paired with natural gas generation. Moderate deployment of these technologies can both lower the cost of decarbonization and greatly diminish the value of added transmission, even if these emerging technologies are costly.
Our work contributes to the literature by evaluating transmission value in least-cost, zero-emissions, and socially-optimal scenarios, while building from existing infrastructure, including transmission. It accounts for how a broad range of technologies, including renewables, batteries, hydrogen, nuclear, and CCS, can substitute for transmission using a model with high temporal and spatial resolution of weather and associated supply and demand.
@@
Our main finding is that an optimal nationwide zero-emissions plan would more than triple current interregional transmission, which validates the conventional wisdom that much more is needed. However, we find that such a system would cost only 4% less than a well-designed zero-emissions system with no new interregional transmission, as a complex array of substitutes from storage to generation mix adjustments can limit excess costs.
@@ 2 Current and idealized future systems
The value of expanding transmission depends on what already exists, including generation and transmission infrastructure. It also depends on the resource availability to build and connect new generation and on expected future demand. The existing grid evolved from primarily local power provision to increasingly interconnected regions. At first, transmission was used mainly to transport power from remote generating plants to urban customers (Decker, 2021). Over time, it became clear that efficiency would be enhanced by connecting population centers to exploit economies of scale with large “base load" power plants that could operate continuously at full capacity (e.g., coal and nuclear), paired with smaller local ramping and peaking power plants that could accommodate time-varying demand (Masters, 2013). A larger, more connected grid also has more inertia and stability in the face of demand shocks, severe weather events, plant outages, and, increasingly, renewable energy intermittency. As a result of these economic forces, regions are more interconnected today than a century ago. Still, the U.S. grid remains highly fragmented, with three nearly disjoint sections, the Eastern, Western, and ERCOT interconnections (panel b of Figure 1), and varying degrees of connectivity between regions within each of these interconnects.
@@
The existing transmission system was not designed for the future, which will likely employ much more solar, wind, batteries, and other forms of storage, given their low and declining costs and lack of CO and other pollution emissions. However, long-duration storage remains expensive, so there is an ongoing debate about what the future generation mix should look like (Jacobson et al., 2018; Pursiheimo, Holttinen and Koljonen, 2019; Qazi, 2022), and how the degree of interregional transmission expansion could factor into the generation mix (Staadecker et al., 2023), its geographic disposition, and overall cost. The future energy system is also policy-dependent. It will hinge on subsidies for clean energy, taxes or other restrictions on carbon dioxide and other pollutant emissions, local market rules and structures, and decisions by state and federal regulators.
@@
To obtain a sense of what an idealized future system could look like, we used Switch (Fripp, 2012; Johnston et al., 2019; Staadecker et al., 2023; Hidalgo-Gonzalez, Johnston and Kammen, 2021; Sánchez-Pérez et al., 2022), a state-of-the-art, open-source power-system planning model, paired with PowerGenome (https://github.com/PowerGenome/PowerGenome), an open-source platform that compiles comprehensive weather and utility system data, to develop a least-cost plan for a zero-carbon power system for the continental United States in 2050. The model co-optimizes potential new generation capacities, storage, interregional transmission, and hourly chronological operation of the system for all 8,760 hours, which is essential for finding true least-cost high-renewable systems given their weather-driven variability and critical dependence on storage and transmission. Technology and fuel cost assumptions are derived mainly from the National Renewable Energy Laboratory’s Annual Technology Baseline (National Renewable Energy Laboratory, 2022). This idealized system of the future, portrayed in panel c of Figure 1, is dominated by solar and wind generation, possesses ample storage in the form of batteries and green hydrogen, and increases interregional transmission capacity by over 200%.
@@
This idealized, zero-emissions system is 68% more costly per MWh than the least-direct-cost system that excludes any account of external CO emissions costs. At a markedly lower cost (45% above least-cost, excluding emissions’ costs), we could achieve roughly 95% decarbonization, a level of decarbonization consistent with the Environmental Protection Agency’s estimated social cost of carbon dioxide of $190 per ton. The difference in cost and emissions between the socially optimal and zero-emissions scenarios is mainly derived from natural gas with carbon capture and storage (CCS) replacing hydrogen. Note that these costs pertain only to generation and transmission and do not account for local distribution costs, which can comprise almost half of the retail electricity costs for residential customers today (Borenstein, Fowlie and Sallee, 2021). Hence, these decarbonization costs are a considerably smaller share of retail prices. However, these idealized systems optimize interregional transmission, roughly tripling it, including building new connections between the three interconnects.
@@ 3 Scenarios considered
Each scenario jointly optimizes generation and storage capacities, transmission, and hourly chronological operation of the system over 52 weeks, subject to emissions and transmission constraints of the given scenario. We use 942 clusters of candidate renewable energy projects, corresponding to approximately 21 TW of potential capacity, with each cluster’s variable capacity factors derived from weather data synchronized with demand. Both weather and demand are based on 2012; details are provided in methods.
To show how interregional transmission influences the cost of decarbonization, we built a series of scenarios with different restrictions on emissions and transmission expansion. These include three emissions scenarios: (i) a least-direct-cost system that ignores external costs of CO and other pollution emissions; (ii) a socially-optimal system that assumes a CO price of $190 per ton; and (iii) a zero-emissions system with no CO emissions. All three of these scenarios use the same projected demand for 2050, which is 74% greater than today’s demand, roughly accounting for the growth of electric vehicles, partial electrification of heat, and electrification of some industrial activities. For each of these three emissions scenarios, we consider three transmission scenarios: (a) existing transmission in 2022; (b) optimized expansion within interconnects, but no expansion between interconnects; and (c) fully-optimized expansion of transmission between regions both within and between interconnects. Crossing the emissions scenarios and transmission scenarios gives a total of nine scenarios labeled below.
@@
The main figures in this manuscript focus on LE, LO, ZE, and ZO and we report results for the other scenarios in the supplement since they are similar to the cases shown here. In each case, all other aspects of the system are co-optimized, subject to transmission constraints. Existing and candidate generation technologies include solar, wind, batteries, hydrogen, natural gas with and without carbon capture, battery storage, nuclear, pumped-storage hydroelectric power, pre-existing hydroelectricity, and distributed solar. Hydrogen production can be produced endogenously via electrolysis and later used for generation via fuel cell, and thus serves as a form of long-term storage. The model accounts for all systems’ operations and maintenance costs, including existing transmission lines. By optimizing all options in conjunction with different transmission constraints, we see how portfolio adjustments compensate for less-than-ideal expansion of long-distance transmission.
@@ 4 Results
This paper uses average load-weighted, long-run marginal cost as the “wholesale cost" metric (the dual of the load-serving constraint in the optimization model). This roughly corresponds to demand-weighted locational marginal prices in wholesale power markets, assuming markets are perfectly competitive and in long-run equilibrium. This metric includes rents to scarce resources and can thus differ slightly from engineering-based measures of levelized cost. Note, however, that we do not account for reserves, excess capacity payments for obsolete or partially obsolete assets, within-region congestion costs, or any sub-optimal investment or operation of the system. The costs also correspond to a single weather year.
@@
Optimizing transmission for the least-direct-cost system minimally reduces the average wholesale cost of power from 37.4 $/MWh under existing transmission to 37.0 $/MWh, a savings of just 0.3% (Overall LE - Overall LO in Figure 2, panel a). This difference is barely visible in the graph due to its small size. Optimally expanding transmission within interconnects reduces overall cost by 0.2 $/MWh, or about half the savings from fully-optimal transmission. Although the least-direct-cost scenarios do not take into account emissions and assume inexpensive future natural gas following NREL-ATB projections, they still include a nearly five-fold expansion of solar and wind capacity relative to 2022 levels and reduce electricity sector CO emissions by 6.9 to 7.8% relative to 2022. These emission reductions occur despite 74% growth in overall demand and thus result in over 50% reduction in emissions intensity (CO per MWh delivered) compared to 2022. Furthermore, the projected growth in demand embodies substantial growth of electric vehicles and electrification of heat, such that economy-wide emissions reductions would likely be much greater; a precise calculation is beyond the scope of this study.
@@
Optimizing transmission for the zero-emissions system reduces the average wholesale cost of power from 64.8 $/MWh under existing transmission to 62.2 $/MWh, a savings of 4% (Overall ZE - Overall ZO in Figure 2). Optimally expanding transmission within interconnects, but not between them, reduces cost by 3.2% (not shown). The relative savings of expanded transmission are greater in the high-renewable zero-emissions system than in the least-direct-cost system with about one-third the amount of wind and solar. And because the direct costs are far higher, net transmission savings are fairly substantial. As shown in Figure 5, optimal transmission scenarios reduce prices slightly throughout the distribution.
@@
The overall cost of decarbonization, defined as the difference between the average marginal cost of a zero-emissions system and a least-direct-cost system, is 25.2–27.4 $/MWh, or 68–73% of the least-direct-cost system, depending on the transmission scenario. If emissions are valued at $190/tCO, the avoided harm in the zero-carbon case is much greater than the extra direct costs in the power system. We find that full decarbonization has total societal costs—including both direct costs and emission impacts—that are 57%–59% below the least-direct-cost case. If, instead of full decarbonization, societal costs are minimized, direct costs increase by 34–36% relative to the least-direct-cost scenario, electricity sector emissions fall by 95.5 to 95.7% relative to 2022, and societal savings rise to 66%. These estimates do not include economic and emissions savings in transportation, heating, and industrial sectors of the economy that would be substantially electrified (supplement Table A1).
@@
The costs of decarbonization and benefits of optimal transmission vary across regions. In Figure 2 panel a, regions are ordered from least to greatest cost in the scenario with zero emissions and existing transmission (ZE). These costs vary from about $47/MWh in TRE_WEST, which encompasses southwestern Texas, to about $91/MWh in NYCM, which contains New York City and Long Island. It is clear from the map (panel b) that the transmission benefit for decarbonization varies across regions; in a few cases, expanding transmission increases the decarbonization cost (shades of red). The precise calculation for each point is the demand-weighted average hourly marginal cost for each region under each scenario. Because the model optimizes the system’s capital and chronological operation, and capital decisions are nearly continuous at the regional scale, the average marginal cost equals the average cost, including positive or negative economic rents to infra-marginal or stranded generation and transmission. We report on rents below and show how transmission changes their allocation, which may be relevant to the political economy of transmission expansion and of generation plants siting.
@@
To fully decarbonize using only existing transmission, investments in storage must increase by 21%, generation by 9%, and hydrogen electrolysis by 0.6% relative to the fully-optimized transmission scenario (Figure 3, panel a ). If transmission can expand optimally only within interconnects, generation and storage capacities change minimally relative to the fully optimized transmission scenario, while hydrogen is minimally used regardless of transmission. In the socially optimized system, limiting to existing transmission causes storage and generation expenditure to increase by 23% and 5%, while hydrogen is minimally used regardless of transmission (shown in the supplement in Table A1 and Figure A4). In general, the mix of generation is similar regardless of transmission but uses slightly more nuclear and solar and less wind when transmission is constrained (panels c and d of Figure 3). These compensating adjustments in generation and storage expansions are somewhat more expensive than the optimized transmission expansions that they replace (Figure 3, panel a). The adjustments are modest, however, relative to the 15-fold increase in renewables and other capacity increases needed for decarbonization.
@@
In addition to the technologies shown, the model includes options for carbon capture and storage (CCS) paired with natural gas and coal generation facilities. These technologies are only selected in the socially optimal scenarios, with details provided in the supplement (Figure A4).
@@
Transmission expansion will have little influence on price variability, while the extent of decarbonization will significantly increase price variability (Figure 5). In competitive markets, wholesale prices equal marginal cost, the distributions of which we can infer from model’s dual variables associated with each hour’s balancing constraint. We show these distributions as cumulative densities because they tend to have significant mass points at zero, and especially large ones in high-renewable scenarios. In the supplement we use multiple bar plots to show the density functions, which may be more intuitive. We discuss likely real-world departures from competitive markets below.
@@
The marginal-cost distributions for the least-direct-cost scenarios have about 90 percent of MWh between 20 and 40 dollars per MWh. About 2 percent of MWh have a marginal cost of zero or nearly so ( $1/MWh), and about five percent have marginal costs above 40 dollars per MWh. The zero marginal-cost times occur during curtailment events (wind and/or solar energy are discarded). The influence of transmission on the marginal cost distribution is so small that it is difficult to discern in the graph.

### v2

@@ Abstract
Solar and wind power are cost-competitive with fossil fuels, yet their intermittent nature presents challenges. Significant temporal and geographic differences in land, wind, and solar resources suggest that long-distance transmission could be particularly beneficial. Using a detailed, open-source model, we analyze optimal transmission expansion jointly with storage, generation, and hourly operations across the three primary interconnects in the United States. Transmission expansion offers far more benefits in a high-renewable system than in a system with mostly conventional generation. Yet while an optimal nationwide plan would have more than triple current interregional transmission, transmission decreases the cost of a 100% clean system by only 7% compared to a plan that relies solely on current transmission. Expanding capacity only within existing interconnects can achieve most of these savings. Adjustments to energy storage and generation mix can leverage the current interregional transmission infrastructure to build a clean power system at a reasonable cost.
@@ 1 Introduction
Although solar and wind power are increasingly affordable and are likely to be central to global decarbonization efforts, their variability presents a considerable challenge to existing systems, which are primarily designed around easily controllable thermal power plants. Interregional transmission can mitigate these challenges by balancing supply and demand across regions and facilitating the transfer of electricity from areas rich in solar and wind resources to those with fewer renewable options. For instance, the desert Southwest receives more solar radiation than other regions, while the Great Plains have abundant wind resources. These aspects of solar and wind contrast sharply with thermal power plants, which perform nearly the same wherever they are placed. Because much of the existing interregional transmission infrastructure was not built with a high-renewable future in mind, it remains uncertain whether current transmission lines are well-suited for the shifting interregional and intertemporal dynamics of an increasingly renewable-powered grid. At the same time, extreme weather events and reliability impacts (e.g., heat waves and wildfires in California and winter storms in Texas) have increased discussion of the need for strong transmission connections between U.S. grids (botterud2024bridging; doe2023needs).
Despite its potential benefits, expanding transmission capacity is fraught with economic, logistical, environmental, and political obstacles. The construction of new transmission lines is historically contentious due to disputes over siting, financing mechanisms, and cost allocation among stakeholders. While increased interregional trade generally improves overall efficiency, some stakeholders—including existing transmission line owners and locally advantaged generators—may face economic losses due to heightened competition (hausman2024). Additionally, transmission projects often traverse regions that do not directly benefit from them, and these communities may understandably resist their siting—particularly when such projects pose environmental and biodiversity risks (biasotto2018power; marshall2002strategic; soderman2006treatment; hyde2018transmission). Furthermore, ERCOT, the isolated Texas grid, is particularly resistant to interconnections due to its unique market structure and longstanding aversion to Federal Energy Regulatory Commission (FERC) oversight. These challenges raise doubts about whether interregional transmission expansion will progress in line with conventional assessments of its benefits and costs.
Previous research has highlighted the key role of transmission expansion in decarbonization (pacala2021; moch2022; joskow2020; davis2023) and explored various expansion strategies (rosellon2003different; olmos2018transmission). However, relatively few studies have provided quantitative estimates of the economic value of transmission expansion in achieving a zero-emission electricity system. To the best of our knowledge, the study by brown2021value is the only work to analyze a 100% renewable future using a capacity expansion model. Their results suggest that upgrading interregional transmission infrastructure could reduce the cost of decarbonized power by approximately 30%, lowering the average electricity price from over $100 per MWh to about $70 per MWh. However, a key limitation of their analysis is its focus on an idealized “greenfield” transmission system, with little attention to existing infrastructure. Moreover, while their model features high temporal resolution, it does not fully account for seasonal balancing solutions, such as green hydrogen or carbon capture and storage (CCS) paired with natural gas. The potential deployment of these technologies—despite their costs—could significantly alter decarbonization pathways and reduce the economic value of additional transmission capacity.
In summary, while prior studies generally agree that transmission expansion is critical for deep decarbonization, there remains a limited body of work that directly quantifies its economic value. The findings of brown2021value reinforce the widespread view that transmission investment is beneficial, but more research is needed to evaluate transmission expansion in a broader range of scenarios, including those that integrate alternative flexibility solutions.
Our study advances this literature by adopting a different modeling approach that evaluates the economic value of transmission in multiple scenarios while explicitly incorporating existing transmission infrastructure. Our model captures how various technologies—including renewables, batteries, hydrogen, nuclear, and CCS—can complement or substitute for transmission expansion. Using high temporal and spatial resolution, our framework reflects weather-driven variability in electricity supply and demand, providing a more comprehensive assessment of the role of transmission in a decarbonized energy system.
@@
Our main finding is that an optimal nationwide zero-emissions plan would more than triple current interregional transmission, which validates the conventional wisdom that much more is needed. However, such a system would cost only 7% less than a well-designed zero-emissions system with no new interregional transmission, as a complex array of substitutes from storage to generation mix adjustments can limit excess costs. Transmission value is less in least-cost and socially optimal scenarios with positive but greatly diminished emissions intensities. We test the robustness of these findings using a proxy to represent extreme weather by boosting demand by 25% on the three consecutive days in each region that are the most costly to balance, which raises transmission value to 7.7%. We also find that transmission expansion increases rents for some resources and regions and decreases them for others. In general, transmission expansion tends to favor regions that have high-quality wind or solar resources.
Our work contributes to the literature in two ways. First, we characterize and evaluate substitutes for optimal transmission expansion in high-renewable systems. We proceed by evaluating transmission value in least-cost, zero-emissions, and socially optimal scenarios while building from existing infrastructure, including transmission. It accounts for how a broad range of technologies, including renewables, batteries, hydrogen, nuclear, and CCS, can substitute for transmission using a model with high temporal and spatial resolution of weather and associated supply and demand. Second, our work evaluates the gains and losses to owners of existing transmission and other generation resources if transmission is expanded optimally.
@@ 2 Current and idealized future systems
The value of expanding transmission depends on what already exists, including generation and transmission infrastructure. It also depends on the resource availability to build and connect new generation and on expected future demand. The existing grid evolved from primarily local power provision to increasingly interconnected regions. At first, transmission was used mainly to transport power from remote generating plants to urban customers (DECKER2021193). Over time, it became clear that efficiency would be enhanced by connecting population centers to exploit economies of scale with large “base load" power plants that could operate continuously at full capacity (e.g., coal and nuclear), paired with smaller local ramping and peaking power plants that could accommodate time-varying demand (masters2013renewable). A larger, more connected grid also has more inertia and stability in the face of demand shocks, severe weather events, plant outages, and, increasingly, renewable energy intermittency. As a result of these economic forces, regions are more interconnected today than a century ago. Still, the U.S. grid remains highly fragmented, with three nearly disjoint sections, the Eastern, Western, and ERCOT interconnections (panel b of Figure 1), and varying degrees of connectivity between regions within each of these interconnects.
@@
The existing transmission system was not designed for the future, which will likely employ much more solar, wind, batteries, and other forms of storage, given their low and declining costs and lack of CO2 and other pollution emissions. However, long-duration storage remains expensive, so there is an ongoing debate about what the future generation mix should look like (jacobson2018matching; pursiheimo2019inter; qazi2022future), and how the degree of interregional transmission expansion could factor into the generation mix (staadecker_2023), its geographic disposition, and overall cost. The future energy system is also policy-dependent. It will hinge on subsidies for clean energy, taxes or other restrictions on carbon dioxide and other pollutant emissions, local market rules and structures, and decisions by state and federal regulators.
@@
To obtain a sense of what an idealized future system could look like, we used Switch (fripp2012switch; johnston2019switch; staadecker_2023; hidalgo2021cost; sanchez2022effect), a state-of-the-art, open-source power-system planning model, paired with PowerGenome (https://github.com/PowerGenome/PowerGenome), an open-source platform that compiles comprehensive weather and utility system data, to develop a least-cost plan for a zero-carbon power system for the continental United States in 2050. The model co-optimizes potential new generation capacities, storage, interregional transmission, and hourly chronological operation of the system for all 8,760 hours, which is essential for finding true least-cost high-renewable systems given their weather-driven variability and critical dependence on storage and transmission. Technology and fuel cost assumptions are derived mainly from the National Renewable Energy Laboratory’s Annual Technology Baseline (nrelatb2022). This idealized system of the future, portrayed in panel c of Figure 1, is dominated by solar and wind generation, possesses ample storage in the form of batteries and green hydrogen and increases interregional transmission capacity by over 200%.
@@
This idealized, zero-emissions system is 78% more costly per MWh than the least-direct-cost system that excludes any account of external CO2 emissions costs. At a markedly lower cost (36% above least-cost, excluding emissions’ costs), we could achieve roughly 95% decarbonization, a level of decarbonization consistent with the Environmental Protection Agency’s estimated social cost of carbon dioxide of $190 per ton. The difference in cost and emissions between the socially optimal and zero-emissions scenarios is mainly derived from natural gas with carbon capture and storage (CCS) replacing hydrogen. Note that these costs pertain only to generation and transmission and do not account for local distribution costs, which can comprise almost half of the retail electricity costs for residential customers today (borenstein2021designing). Hence, these decarbonization costs are a considerably smaller share of retail prices. However, these idealized systems optimize interregional transmission, more than tripling it, including building new connections between the currently almost disjoint three interconnects (Eastern, Western, and ERCOT).
@@ 3 Scenarios considered
We consider a range of scenarios to evaluate the value of transmission relative to potential substitutes. Each scenario jointly optimizes generation and storage capacities, transmission, and hourly chronological operation of the system for each of 52 weeks, subject to emissions and transmission constraints of the given scenario. We use 1014 clusters of candidate renewable energy projects, corresponding to approximately 27 TW of potential capacity, with each cluster’s variable capacity factors derived from weather data synchronized with demand. Both weather and demand are based on actual values from 2012; details are provided in methods.
To show how interregional transmission influences the cost of decarbonization, we built a series of scenarios with different restrictions on emissions and transmission expansion. These include three emissions scenarios: (i) a least-direct-cost system that ignores external costs of CO2 and other pollution emissions; (ii) a socially-optimal system that assumes a CO2 price of $190 per ton; and (iii) a zero-emissions system with no CO2 emissions. All three of these scenarios use the same projected demand for 2050, which is 74% greater than today’s demand, roughly accounting for the growth of electric vehicles, partial electrification of heat, and some industrial activities. For each of these three emissions scenarios, we consider three transmission scenarios: (a) existing transmission in 2022; (b) optimized expansion within interconnects, but no expansion between interconnects; and (c) fully-optimized expansion of transmission between regions both within and between interconnects. Crossing the emissions scenarios and transmission scenarios gives a total of nine scenarios labeled below.
@@
In addition, we consider three supplemental scenarios:
BE | Zero-emissions, existing trans, boosted demand. |
BO | Zero-emissions, fully-optimal trans, boosted demand. |
ZE+ | Zero-emissions, with transmission expansion limited to 25% of existing. |
The BE and BO scenarios are like ZE and ZO except we increase demand for all hours by 25 percent on the three consecutive days that are most difficult to serve in each region in the ZE scenario. These scenarios test our conclusions against extreme conditions that are likely to favor transmission. This diagnostic may exaggerate the value of transmission because the difficult days are determined and boosted idiosyncratically, where in reality unusual events are likely to be correlated across regions, which would make transmission less valuable. The ZE+ scenario considers the value of increasing transmission modestly, and sheds light on the most valuable expansions possible.
The main figures in this manuscript focus on LE, LO, ZE, and ZO, and we report results for the other scenarios in the supplement since they are similar to the cases shown here. In each case, all other aspects of the system are co-optimized, subject to transmission constraints. Existing and candidate generation technologies include solar, wind, batteries, hydrogen, natural gas with and without carbon capture, battery storage, nuclear, pumped-storage hydroelectric power, preexisting hydroelectricity, and distributed solar. Hydrogen production is produced endogenously via electrolysis and later used for generation via fuel cells, and thus serves as a form of long-term storage. The model accounts for all systems’ operations and maintenance costs, including existing transmission lines. By optimizing all options in conjunction with different transmission constraints, we see how portfolio adjustments compensate for less-than-ideal expansion of long-distance transmission.
@@ 4 Results
This paper uses average load-weighted, long-run marginal cost as the “wholesale cost" metric (the dual of the load-serving constraint in the optimization model). This roughly corresponds to demand-weighted locational marginal prices (LMP) in wholesale power markets, assuming markets are perfectly competitive and in long-run equilibrium. This metric includes rents to scarce resources and can thus differ slightly from engineering-based measures of levelized cost. We do not account for excess capacity payments, costs of obsolete or partially obsolete assets, within-region congestion costs, or any sub-optimal investment or operation costs of the system.
@@
Optimizing transmission for the least-direct-cost system minimally reduces the average wholesale cost of power from 37.9 $/MWh under existing transmission to 37.6 $/MWh, a savings of just 0.8% (Overall LE - Overall LO in Figure 2, panel a). This difference is barely visible in the graph due to its small size. Optimally expanding transmission only within interconnects reduces overall cost by 0.28 $/MWh, or almost all the savings from fully-optimal transmission. Although the least-direct-cost scenarios do not take into account emissions and assume inexpensive future natural gas following NREL-ATB projections, they still include a nearly five-fold expansion of solar and wind capacity relative to 2022 levels and increase electricity sector CO2 emissions by around 10% relative to 2022. This emission increase follows from a 74% growth in overall demand and thus results in a 30% decrease in emissions intensity (CO2 per MWh delivered) compared to 2022. Furthermore, the projected growth in demand embodies substantial growth of electric vehicles and electrification of heat, such that economy-wide emissions reductions would likely be much greater; a precise calculation is beyond the scope of this study.
@@
Optimizing transmission for the zero-emissions system reduces the average wholesale cost of power from 70.4 $/MWh under existing transmission to 65.5 $/MWh, a savings of 7.0% (Overall ZE - Overall ZO in Figure 2). Optimally expanding transmission within interconnects reduces cost by 6.1% (not shown in the Figure). The relative savings of expanded transmission are thus much greater in the high-renewable zero-emissions system than in the least-direct-cost system. As shown in Figure 3, optimal transmission scenarios reduce wholesale marginal costs slightly throughout the distribution (also see Figure A3, in the supplement). If we limit transmission expansion to 25% more than existing, wholesale power costs equal 69.3 $/MWh, achieving about one 1/5 the benefits of optimal transmission with about 1/12 the expansion (Table A1). This exercise also highlights the most critical expansions, which we illustrate in the supplement (Figure A9). These expansions key expansions strengthen ties between ERCOT and the Eastern interconnect, the East and West interconnects, and Northeastern connections, especially around New York City.
@@
The overall cost of decarbonization, defined as the difference between the average LMP of a zero-emissions system and a least-direct-cost system, is 27.9–32.5 $/MWh, or 74–86% of the least-direct-cost system, depending on the transmission scenario. If emissions are valued at $190/tCO2, the avoided harm in the zero-carbon case is much greater than the extra direct costs in the power system. We find that full decarbonization has total societal costs—including both direct costs and emission impacts—that are 57%–59% below the least-direct-cost case. If societal costs are minimized instead of full decarbonization, direct costs increase by 22–28% (supplement Table A1).
@@
The costs of decarbonization and benefits of optimal transmission vary across regions. In Figure 2 panel a, regions are ordered from least to greatest cost in the scenario with zero emissions and existing transmission (ZE). These costs vary from about $50/MWh in TREW, which encompasses western Texas, to about $120/MWh in NYCM, which contains New York City and Long Island. It is clear from the map (panel b) that the transmission benefit for decarbonization varies across regions; in one case, expanding transmission increases the decarbonization cost (NYUP). The precise calculation for each point is the demand-weighted average hourly marginal cost for each region under each scenario. A neighboring region (NYCW), which encompasses New York City and Long Island, benefits the most from transmission expansion, reducing LMP by approximately 25%. Because the model optimizes the system’s capital and chronological operation, and capital decisions are nearly continuous at the regional scale, the average marginal cost equals the average cost, including positive or negative economic rents to infra-marginal or sunk generation and transmission assets. We report on rents below and show how transmission changes their allocation, which may be relevant to the political economy of transmission expansion and of siting generation plants.
@@
To fully decarbonize using only existing transmission, investments in storage must increase by 32%, generation by 14%, and hydrogen electrolysis by 1.4% relative to the fully-optimized transmission scenario (Figure 4, panel a). If transmission can expand optimally only within interconnects, generation and storage capacities change minimally relative to the fully optimized transmission scenario, while a small amount of hydrogen is used regardless of transmission. In the socially optimized system, limiting to existing transmission requires storage and generation expenditure to increase by 29% and 22%, again with limited hydrogen use regardless of transmission (shown in the supplement in Table A1 and Figure A4). In general, the mix of generation is similar regardless of transmission but uses slightly more nuclear and solar and less wind when transmission is constrained (panels c and d of Figure 4). These compensating adjustments in generation and storage expansions are somewhat more expensive than the optimized transmission expansions that they replace (Figure 4, panel a). The adjustments are modest, however, relative to the eight- to nine-fold increase in solar and wind (from the end of 2023), among other capacity increases, needed for decarbonization.
@@
In addition to the technologies shown, the model includes options for carbon capture and storage (CCS) paired with natural gas and coal generation facilities. Because CCS is not zero-emissions, it is only selected in the socially optimal scenarios, with details provided in the supplement (Figure A4).
When we make the most costly days in each region even more difficult to serve by boosting demand 25 percent — scenarios BE, BO, we find the value of transmission increases from 7% to 7.6%, which indicates our findings are unlikely the result of limited weather and demand outcomes (Figure A8).
@@
Transmission expansion has little influence on price variability, while the extent of decarbonization significantly increases price variability (Figure 3). In competitive markets, wholesale prices equal marginal cost, the distributions of which we can infer from model’s dual variables associated with each hour’s balancing constraint. We show these distributions as bar plots with binned ranges because of the significant mass point at zero and to ease comparison of scenarios with and without transmission expansion. The supplement shows cumulative density functions (Figure 2). We discuss possible real-world departures from competitive pricing below.
@@
The marginal-cost distributions for the least-direct-cost scenarios have about 90 percent of MWh between 20 and 40 dollars per MWh. About 2 percent of MWh have a marginal cost of zero or nearly so ( $1/MWh), and about 4 percent have marginal costs above 40 dollars per MWh. The zero marginal-cost times occur during curtailment events (wind and/or solar energy are discarded).